The length of time it takes college students to find a parking spot in the library parking lot follows a
normal distribution with a mean of 5.5 minutes and a standard deviation of 1 minute. Find the
probability that a randomly selected college student will take between 4.0 and 6.5 minutes to find a
parking spot in the library lot.
Answer #1
4-5.5 = -1.5
z =-1.5/1 = -1.5 or 1.5sigma below mean

6.5 - 5.5 = 1
z = 1/1 = 1 sigma above mean
so what fraction of a normal distribution lies between z = -1.5 and z = + 1?
from table
F(-1.5) = .067 are below -1.5
F(1) = .841 are below +1
so .842-.067 are between
answered by: Shawn11